### v1

@@ 2.2 FlexGT and Acc-FlexGT Algorithms
0: Initial points $z_{i,0}=x_{i,0}\in\mathbb{R}^{p}$ and $y_{i,0}=\nabla_{x}f_{i}\left(x_{i,0};\xi_{i,0}\right)$, number of communication and computation steps $\alpha,\beta\geqslant 1$, and stepsize $\gamma>0$.
1:for round $k=0,1,\dots$, each node $i\in[n]$do
@@
4:   Obtain an unbiased noisy gradient $\nabla f_{i}\left(z_{i,l+1};\xi_{i,l+1}\right)$ by querying $\mathcal{S}\mathcal{O}$.
5:   Perform local updates:
$x_{i,l+1}$ $=x_{i,l}-\gamma y_{i,l}\quad\quad\text{\# gradient descent;}$
$z_{i,l+1}$ $=z_{i,l}\quad\quad\text{\# snapshot point;}$
$y_{i,l+1}$ $=y_{i,l}+\nabla f_{i}\left(z_{i,l+1};\xi_{i,l+1}\right)-\nabla f_{i}\left(z_{i,l};\xi_{i,l}\right)\quad\text{\# gradient tracking}$
6:  end for
7:  Let $\theta_{i,0}=\theta_{i,-1}=\left\{x_{i,\beta\left(k+1\right)},y_{i,\beta\left(k+1\right)}\right\}$.
8:  for $s=0,1,\dots,\alpha-1$ do
9:   FlexGT: Direct inter-node communication:
$\vskip-8.5359pt\theta_{i,s+1}=\sum_{j\in\mathcal{N}_{i}}{w_{ij}\theta_{j,s}}.$
10:   Acc-FlexGT: Accelerated gossip communication:
$\vskip-2.84544pt\theta_{i,s+1}=\left(1+\eta\right)\sum_{j\in\mathcal{N}_{i}}{w_{ij}\theta_{j,s}-\eta}\theta_{i,s-1},$
@@
11:  end for
12:end for
Figure 1: Flow diagram illustrating the computation and communication of FlexGT. The variable $\psi_{i,l}$, $i=1,\dots,n$, represents the collection of $x_{i,l}$, $y_{i,l}$, and $z_{i,l}$. Only $x_{i,l}$ and $y_{i,l}$ are communicated. The communication is weighted by the matrix $W$.
We introduce the following notations to represent the collection of decision parameters, GT variables, and snapshot points, respectively:
@@
(3a) $\mathbf{x}_{\beta\left(k+1\right)}$ $=\bar{W}\left(\mathbf{x}_{\beta k}-\gamma\sum_{j=0}^{\beta-1}{\mathbf{y}_{\beta k+j}}\right),$
(3b) $\mathbf{y}_{\beta\left(k+1\right)}$ $=\bar{W}\left(\mathbf{y}_{\beta k}+\nabla G_{\beta\left(k+1\right)}-\nabla G_{\beta k}\right),$
@@
$\nabla G_{t}:=\left[\dots,\nabla f_{i}\left(z_{i,t};\xi_{i,t}\right),\dots\right]^{\top}\in\mathbb{R}^{n\times p}$
@@
$\nabla F_{t}:=\mathbb{E}\left[\nabla G_{t}\right]=\left[\dots,\nabla f_{i}\left(z_{i,t}\right),\dots\right]^{\top}\in\mathbb{R}^{n\times p}.$
We note that $\bar{W}$ takes different forms depending on the communication protocol employed. In the case of FlexGT, we have $\bar{W}=W^{\alpha}$. Otherwise, when the accelerated gossip communication protocol is used as in Acc-FlexGT, we have $\bar{W}=M_{\alpha}$ with $M_{\alpha}$ obtained through $\alpha$ communication steps:
@@
with $M_{-1}=M_{0}=I$. Although in both cases there are $\alpha$ communications per round, the difference is in prior knowledge, e.g., the spectral gap of the graph, which leads to distinct communication properties, as described in the following lemma.
Lemma 1.
@@
(4) $\bar{\rho}_{W}=\left\|W^{\alpha}-\mathbf{J}\right\|^{2}\leqslant\rho_{W}^{\alpha};$
@@
(5) $\bar{\rho}_{W}=\left\|M_{\alpha}-\mathbf{J}\right\|^{2}\leqslant 2\left(1-\sqrt{1-\sqrt{\rho_{W}}}\right)^{2\alpha}.$
Equation (4) is straightforward to derive based on Assumption 4, and (5) was stated at Proposition 3 in [28].
@@
(6) ${\mathbf{z}_{\beta k+t}}=\,\,\mathbf{x}_{\beta k},\quad t=0,1,\dots,\beta-1,$
which implies that $\mathbf{y}$ tracks the stochastic gradient at the snapshot point $\mathbf{z}$ instead of the current decision variable $\mathbf{x}$, i.e.,
(7) $\mathbf{1}^{\top}{\mathbf{y}_{\beta k+t}}=\mathbf{1}^{\top}{\nabla G_{\beta k+t}}=\sum_{i=1}^{n}{\nabla f_{i}\left(x_{i,\beta k};\xi_{i,\beta k+t}\right)}.$
This prevents the tracking variables from diverging excessively due to multiple local updates, i.e., the gradient difference $\nabla G_{t+1}-\nabla G_{t}$ remains controlled, which is crucial for stability in the presence of heterogeneous data. This method differs from the existing snapshot GT approach [43, 10], which randomly skips updating the tracking variables while performing their consensus at each iteration, potentially leaving computation resources idle. In contrast, FlexGT utilizes the stochastic gradient at each iteration to slightly update the tracking variables and perform communication periodically to avoid extra costs from redundant initialization.
@@ 3 Main Results
(8) $V_{t}=\left\|\bar{x}_{t}-x^{*}\right\|^{2}+c_{x}\left\|\tilde{\mathbf{x}}_{t}\right\|^{2}+c_{y}\left\|\tilde{\mathbf{y}}_{t}\right\|^{2},$
where $x^{*}$ indicates the global optimal solution of problem (1), $t$ denotes the iteration, $c_{x}$ and $c_{y}$ are coefficients to be properly designed later, and
$\tilde{\mathbf{x}}_{t}=\mathbf{x}_{t}-\mathbf{1}\bar{x}_{t},\quad\bar{x}_{t}=\mathbf{1}^{\top}\mathbf{x}_{t}/n,\quad\tilde{\mathbf{y}}_{t}=\mathbf{y}_{t}-\mathbf{1}\bar{y}_{t},\quad\bar{y}_{t}=\mathbf{1}^{\top}\mathbf{y}_{t}/n.$
Table 2: Communication and computation complexity of FlexGT for strongly convex, convex, and nonconvex objective functions.
@@
FlexGT Co $\frac{\beta L}{\left(1-\rho_{W}^{\alpha}\right)^{2}\epsilon}+\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\beta\sqrt{L\rho_{W}^{\alpha}\sigma^{2}}}{\left(1-\rho_{W}^{\alpha}\right)^{3/2}\epsilon^{3/2}}$ $\frac{\alpha L}{\left(1-\rho_{W}^{\alpha}\right)^{2}\epsilon}+\frac{\alpha\sigma^{2}}{n\beta\epsilon^{2}}+\frac{\alpha\sqrt{L\rho_{W}^{\alpha}\sigma^{2}}}{\left(1-\rho_{W}^{\alpha}\right)^{3/2}\epsilon^{3/2}}$
@@
Table 3: Communication and computation complexity of Acc-FlexGT for strongly convex, convex, and nonconvex objective functions. The last column presents the corresponding best-known or optimal iteration complexity.
@@
SC $\beta\frac{L}{\mu}+\frac{\sigma^{2}}{n\mu^{2}\epsilon}$ $\left(\frac{\sigma^{2}}{n\beta\mu^{2}\epsilon}+\frac{L}{\mu}\right)\frac{1}{\sqrt{1-\sqrt{\rho_{W}}}}$ $\frac{\sigma^{2}}{n\mu^{2}\epsilon}+\frac{\sqrt{L\sigma^{2}}}{\mu\sqrt{1-\sqrt{\rho_{W}}}\epsilon^{1/2}}$ [10] a
Acc- Co $\frac{L\sigma^{2}}{n\epsilon^{2}}+\frac{\beta\sqrt{L}}{\epsilon}$ $\left(\frac{L\sigma^{2}}{n\beta\epsilon^{2}}+\frac{\sqrt{L}}{\epsilon}\right)\frac{1}{\sqrt{1-\sqrt{\rho_{W}}}}$ / b
@@
    The mark "/" indicates no relevant result in the existing literature.
@@ 3.1 Convergence
(9) $\gamma=\mathcal{O}\left(\min\left\{1,\frac{\left(1-\bar{\rho}_{W}\right)^{2}}{\bar{\rho}_{W}}\right\}\right)\frac{1}{\beta L}$
@@
(10) $\mathbb{E}\left[V_{\beta\left(k+1\right)}\right]$ $\leqslant\underset{\mathrm{linear\,\,rate}}{\underbrace{\left(1-\min\left\{\frac{\mu\beta\gamma}{2},\frac{1-\bar{\rho}_{W}}{8}\right\}\right)}}\mathbb{E}\left[V_{\beta k}\right]$
@@
(11) $\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)-f\left(x^{*}\right)\right]}$ $\leqslant\underset{\mathrm{sublinear\,\,rate}}{\underbrace{\frac{2\mathbb{E}\left[V_{0}\right]}{\gamma\beta K}}}+\underset{\mathrm{centralized\,\,error}\mathrm{}}{\underbrace{\frac{2\gamma\sigma^{2}}{n}}}$
$\quad+\underset{\mathrm{topology\,\,effect}}{\underbrace{\frac{3328\gamma^{2}\beta^{2}L\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{3}}\sigma^{2}}}.$
@@
(12) $\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]}$ $\leqslant\underset{\mathrm{sublinear\,\,rate}}{\underbrace{\frac{8\mathbb{E}\left[f\left(\bar{x}_{0}\right)-f^{*}\right]}{\gamma\beta K}+\frac{8\gamma^{2}\beta^{2}L^{2}\bar{\rho}_{W}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{0}\right\|^{2}\right]}{n\left(1-\bar{\rho}_{W}\right)^{3}K}}}$
@@
and does not require the assumption of uniformly bounded stochastic gradients [29], making it more robust in non-i.i.d. settings. Compared to the best-known results for DSGT ($\alpha=\beta=1$) without acceleration in [18], our findings recover their results in all three settings  [^1] Under Assumption 4 on the graph and using the same convergence metric as in this work. and extend to more general cases with $\alpha\geqslant 1$ and $\beta\geqslant 1$. Note that K-GT [30] is a special case of FlexGT with $\alpha=1$ and $\beta>1$, and the communication complexity obtained in this work improves upon that of K-GT by a factor of $1/\sqrt{1-\rho_{W}}$ in terms of the topology effect for the nonconvex setting, while K-GT experiences lower stochastic gradient noise.
Based on the convergence results in Theorem 1, we can further derive the iteration complexity of FlexGT for different objective functions as follows.
Corollary 1.
Suppose Assumptions 24 hold. Let the stepsize satisfy the condition (9) given in Theorem 1, and denote by $\hat{K}$ the number of iterations required for FlexGT to reach an arbitrary accuracy $\epsilon>0$.
@@
(13) $\hat{K}={\mathcal{O}}\left(\frac{L}{\left(1-\bar{\rho}_{W}\right)^{2}\mu}\log\frac{\left[V_{0}\right]}{\epsilon}+\frac{\sigma^{2}}{\beta\mu^{2}n\epsilon}+\frac{\sqrt{L\bar{\rho}_{W}\sigma^{2}}}{\mu^{3/2}\left(1-\bar{\rho}_{W}\right)^{3/2}\epsilon^{1/2}}\right).$
@@
(14) $\hat{K}=\mathcal{O}\left(\frac{L}{\left(1-\bar{\rho}_{W}\right)^{2}\epsilon}+\frac{\sigma^{2}}{\beta n\epsilon^{2}}+\frac{\sqrt{L\bar{\rho}_{W}\sigma^{2}}}{\left(1-\bar{\rho}_{W}\right)^{3/2}\epsilon^{3/2}}\right)\mathbb{E}\left[V_{0}\right].$
@@
(15) $\hat{K}=\mathcal{O}$ $\left(\frac{L}{\left(1-\bar{\rho}_{W}\right)^{2}\epsilon}+\frac{R_{0}}{\epsilon}+\frac{L\sigma^{2}}{\beta n\epsilon^{2}}\right.\left.+\frac{L\sqrt{\bar{\rho}_{W}\sigma^{2}}}{\left(1-\bar{\rho}_{W}\right)^{3/2}\epsilon^{3/2}}\right)\left(\mathbb{E}\left[f\left(\bar{x}_{0}\right)-f^{*}\right]\right),$
@@
Proof.
Due to space constraints, we present only the main ideas of the proof while omitting the detailed steps. Denote by $\hat{\gamma}$ the obtained upper bound of the stepsize in (9) and
the constant part of the centralized error and the topology effect part in (10)–(12) controlled by the stepsize $\gamma$, respectively, where $\upsilon_{0}$ denotes either $\mathbb{E}\left[V_{0}\right]$ or $\mathbb{E}\left[f\left(\bar{x}_{0}\right)-f^{*}\right]$, depending on the specific case under consideration. Then, we choose a sufficiently small $\gamma$ such that the centralized error and topology effect part have the same order as the linear or sublinear part in terms of the number of iterations $\hat{K}$. To this end, letting the stepsize $\gamma$ satisfy
$\min\left\{\frac{\ln\left(\max\left\{2,\beta^{2}\mu^{2}\hat{K}/\varUpsilon_{1},\beta^{3}\mu^{3}\hat{K}^{2}/\varUpsilon_{2}\right\}\right)}{\beta\mu\hat{K}},\,\,\hat{\gamma}\right\}$
for the strongly convex case; and
$\min\left\{\frac{1}{\sqrt{\varUpsilon_{1}\hat{K}}},\,\,\frac{1}{\sqrt[3]{\varUpsilon_{2}\hat{K}}},\,\,\hat{\gamma}\right\}$
for convex and nonconvex cases, we obtain the complexity results. Similar techniques can be found in [45, 19, 13].
Based on the iteration complexity derived in Corollary 1 and the number of local updates and communications per iteration, we can derive the communication and computation complexity of FlexGT for different objective functions, as summarized in Table 2. To the best of our knowledge, FlexGT achieves the best-known complexity among GT-based methods for solving problem in most settings (cf. Remark 1).
Remark 2.
Table 2 highlights the impact of $\alpha$ and $\beta$ on the communication and computation complexity of FlexGT. Unlike previous work, our results reveal an inherent trade-off between communication and computation. Particularly, as the number of computation steps $\beta$ increases, the communication complexity decreases while the computation complexity increases. Conversely, as the number of communication steps $\alpha$ increases, the computation complexity decreases, while the communication complexity may either increase or decrease depending on $\rho_{W}$. This highlights the need for a careful design of $\alpha$ and $\beta$ to balance communication and computation costs according to the specific application requirements.

### v2

@@ 2.2 FlexGT and Acc-FlexGT Algorithms
0: Initial points $z_{i,0}=x_{i,0}\in\mathbb{R}^{p}$ and $y_{i,0}=\nabla f_{i}\left(x_{i,0};\xi_{i,0}\right)$, number of communication and computation steps $\alpha,\beta\geqslant 1$, and stepsize $\gamma>0$.
1:for round $k=0,1,\dots$, each node $i\in[n]$, do
@@
4:   Update decision variable:
$x_{i,l+1}=x_{i,l}-\gamma y_{i,l}.\vskip-14.22636pt$
5:   Let $z_{i,l+1}=z_{i,l}$.
[…]
7:   Update gradient tracking variable:
$y_{i,l+1}=y_{i,l}+g_{i,l+1}-g_{i,l}.$
8:  end for
9:  Let $\theta_{i,0}=\theta_{i,-1}=\left\{x_{i,\beta\left(k+1\right)},y_{i,\beta\left(k+1\right)}\right\}$.
10:  for $s=0,1,\dots,\alpha-1$ do
11:   FlexGT: Direct inter-node communication:
$\theta_{i,s+1}=\sum_{j\in\mathcal{N}_{i}}{w_{ij}\theta_{j,s}}.$
12:   Acc-FlexGT: Accelerated gossip communication:
$\theta_{i,s+1}=\left(1+\eta\right)\sum_{j\in\mathcal{N}_{i}}{w_{ij}\theta_{j,s}-\eta}\theta_{i,s-1},$
@@
13:  end for
14:end for
We first introduce the notations for the collections of decision variables, GT variables, and snapshot points, respectively:
@@
$\mathbf{x}_{\beta\left(k+1\right)}$ $=\bar{W}\left(\mathbf{x}_{\beta k}-\gamma\sum_{j=0}^{\beta-1}{\mathbf{y}_{\beta k+j}}\right),$ (3a)
$\mathbf{y}_{\beta\left(k+1\right)}$ $=\bar{W}\left(\mathbf{y}_{\beta k}+G_{\beta\left(k+1\right)}-G_{\beta k}\right),$ (3b)
@@
$G_{t}:=\left[\dots,g_{i,t},\dots\right]^{\top}\in\mathbb{R}^{n\times p},\quad g_{i,t}=\nabla f_{i}\left(z_{i,t};\xi_{i,t}\right)$
@@
$\nabla F_{t}:=\mathbb{E}\left[G_{t}\right]=\left[\dots,\nabla f_{i}\left(z_{i,t}\right),\dots\right]^{\top}\in\mathbb{R}^{n\times p}.$
We note that the form of $\bar{W}$ varies with the communication protocol used, as specified in lines 11 and 12 of Algorithm 1. In the case of FlexGT (cf., line 11), we have $\bar{W}=W^{\alpha}$. Otherwise, when the accelerated gossip communication protocol is used as in Acc-FlexGT (cf., line 12), we have $\bar{W}=M_{\alpha}$ with $M_{\alpha}$ obtained through $\alpha$ communication steps:
@@
with $M_{-1}=M_{0}=I$. Note that $\bar{W}$ is still a doubly-stochastic matrix. Although both cases involve $\alpha$ communications per round, they differ in the available prior knowledge, i.e., whether the spectral gap of the graph is known a priori for designing the stepsize $\eta$, which results in distinct communication properties, as stated in the following lemma.
Lemma 1
@@
$\bar{\rho}_{W}=\left\|W^{\alpha}-\mathbf{J}\right\|^{2}\leqslant\rho_{W}^{\alpha};$ (4)
@@
$\bar{\rho}_{W}=\left\|M_{\alpha}-\mathbf{J}\right\|^{2}\leqslant 2\left(1-\sqrt{1-\sqrt{\rho_{W}}}\right)^{2\alpha}.$ (5)
Equation (4) is straightforward to derive based on Assumption 4, and (5) was stated at Proposition 3 in [27]. The difference in $\bar{\rho}_{W}$ reflects the difference in the consensus speed achieved by different communication strategies under the same number of communication rounds.
@@
${\mathbf{z}_{\beta k+t}}=\,\,\mathbf{x}_{\beta k},\quad t=0,1,\dots,\beta-1,$ (6)
which implies that $\mathbf{y}$ tracks the stochastic gradient evaluated at the snapshot point $\mathbf{z}$, rather than at the current decision variable $\mathbf{x}$ as in conventional GT methods [36], i.e.,
$\mathbf{1}^{\top}{\mathbf{y}_{\beta k+t}}=\mathbf{1}^{\top}{G_{\beta k+t}}=\sum_{i=1}^{n}{\nabla f_{i}\left(x_{i,\beta k};\xi_{i,\beta k+t}\right)}.$ (7)
This prevents the tracking variables from diverging excessively due to multiple local updates, i.e., the gradient difference $G_{t+1}-G_{t}$ remains controlled, which is crucial for stability in the presence of heterogeneous data. This method differs from the existing snapshot GT approach [42, 10], which randomly skips updating the tracking variables while performing their consensus at each iteration. In contrast, FlexGT utilizes the stochastic gradient at each iteration to slightly update the tracking variables and perform communication periodically to avoid extra costs from redundant initialization.
@@ 3 Main Results
$V_{t}=\left\|\bar{x}_{t}-x^{*}\right\|^{2}+c_{x}\left\|\tilde{\mathbf{x}}_{t}\right\|^{2}+c_{y}\left\|\tilde{\mathbf{y}}_{t}\right\|^{2},$ (8)
where $x^{*}$ indicates the global optimal solution of Problem (1), $t$ denotes the iteration, $c_{x}$ and $c_{y}$ are coefficients to be properly designed later, and the consensus error and GT error are defined as follows:
$\tilde{\mathbf{x}}_{t}$ $=\mathbf{x}_{t}-\mathbf{1}\bar{x}_{t},\quad\bar{x}_{t}=\mathbf{1}^{\top}\mathbf{x}_{t}/n,$
$\tilde{\mathbf{y}}_{t}$ $=\mathbf{y}_{t}-\mathbf{1}\bar{y}_{t},\quad\bar{y}_{t}=\mathbf{1}^{\top}\mathbf{y}_{t}/n.$
Table 2: Communication and computational complexity of FlexGT required to reach an accuracy level of $\epsilon>0$ for strongly convex, convex, and nonconvex objective functions.
@@
FlexGT Cvx $\frac{\beta L}{\left(1-\rho_{W}^{\alpha}\right)^{2}\epsilon}+\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\beta\sqrt{L\rho_{W}^{\alpha}\sigma^{2}}}{\left(1-\rho_{W}^{\alpha}\right)^{3/2}\epsilon^{3/2}}$ $\frac{\alpha L}{\left(1-\rho_{W}^{\alpha}\right)^{2}\epsilon}+\frac{\alpha\sigma^{2}}{n\beta\epsilon^{2}}+\frac{\alpha\sqrt{L\rho_{W}^{\alpha}\sigma^{2}}}{\left(1-\rho_{W}^{\alpha}\right)^{3/2}\epsilon^{3/2}}$
@@
Table 3: Communication and computational complexity of Acc-FlexGT required to reach an accuracy level of $\epsilon>0$ for strongly convex, convex, and nonconvex objective functions. The last column “Iteration” presents the corresponding best-known or optimal iteration complexity.
@@
SC $\frac{\sigma^{2}}{n\mu^{2}\epsilon}+\frac{\beta L}{\mu}$ $\left(\frac{\sigma^{2}}{n\beta\mu^{2}\epsilon}+\frac{L}{\mu}\right)\frac{1}{\sqrt{1-\sqrt{\rho_{W}}}}$ $\frac{\sigma^{2}}{n\mu^{2}\epsilon}+\frac{\sqrt{L\sigma^{2}}}{\mu\sqrt{1-\sqrt{\rho_{W}}}\epsilon^{1/2}}$ [10] a
Acc- Cvx $\frac{L\sigma^{2}}{n\epsilon^{2}}+\frac{\beta\sqrt{L}}{\epsilon}$ $\left(\frac{L\sigma^{2}}{n\beta\epsilon^{2}}+\frac{\sqrt{L}}{\epsilon}\right)\frac{1}{\sqrt{1-\sqrt{\rho_{W}}}}$ / b
@@
    The mark ”/” indicates no relevant result in the existing literature.
@@ 3.1 Convergence
$\gamma=\mathcal{O}\left(\min\left\{1,\frac{\left(1-\bar{\rho}_{W}\right)^{2}}{\bar{\rho}_{W}}\right\}\right)\frac{1}{\beta L}$ (9)
@@
$\mathbb{E}\left[V_{\beta\left(k+1\right)}\right]$ $\leqslant\underset{\mathrm{linear\,\,rate}}{\underbrace{\left(1-\min\left\{\frac{\mu\beta\gamma}{2},\frac{1-\bar{\rho}_{W}}{8}\right\}\right)}}\mathbb{E}\left[V_{\beta k}\right]$ (10)
@@
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[f\left(\bar{x}_{\beta k}\right)-f\left(x^{*}\right)\right]}$ (11)
$\leqslant\underset{\mathrm{sublinear\,\,rate}}{\underbrace{\frac{2\mathbb{E}\left[V_{0}\right]}{\gamma\beta K}}}+\underset{\mathrm{centralized\,\,error}\mathrm{}}{\underbrace{\frac{2\gamma\sigma^{2}}{n}}}+\underset{\mathrm{topology\,\,effect}}{\underbrace{\frac{3328\gamma^{2}\beta^{2}L\bar{\rho}_{W}}{\left(1-\bar{\rho}_{W}\right)^{3}}\sigma^{2}}}.$
@@
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}\left[\left\|\nabla f\left(\bar{x}_{\beta k}\right)\right\|^{2}\right]}$ (12)
$\leqslant\underset{\mathrm{sublinear\,\,rate}}{\underbrace{\frac{8\mathbb{E}\left[f\left(\bar{x}_{0}\right)-f^{*}\right]}{\gamma\beta K}+\frac{8\gamma^{2}\beta^{2}L^{2}\bar{\rho}_{W}\mathbb{E}\left[\left\|\tilde{\mathbf{y}}_{0}\right\|^{2}\right]}{n\left(1-\bar{\rho}_{W}\right)^{3}K}}}$
@@
and does not require the assumption of uniformly bounded stochastic gradients [28], making it more robust in non-i.i.d. settings. Compared to the best-known results for DSGT ($\alpha=\beta=1$) without acceleration in [19], our findings recover their results in all three settings  [^1] Under Assumption 4 on the graph and using the same convergence metric as in this work. and extend to more general cases with $\alpha\geqslant 1$ and $\beta\geqslant 1$. Compared K-GT [29] in the nonconvex setting, the communication complexity obtained in this work improves upon that of K-GT by a factor of $1/\sqrt{1-\rho_{W}}$ in terms of the topology effect, whereas K-GT benefits from a lower level of stochastic gradient noise.
Based on the convergence results in Theorem 1, we can further derive the iteration complexity of FlexGT for different objective functions in the following corollary.
Corollary 1
Suppose Assumptions 24 hold. Let the stepsize satisfy the condition (9) given in Theorem 1, and denote by $\hat{K}$ the number of rounds required for FlexGT to reach an arbitrary accuracy $\epsilon>0$. Then,
@@
$\hat{K}={\mathcal{O}}\Biggl($ $\frac{L}{\left(1-\bar{\rho}_{W}\right)^{2}\mu}\log\frac{\mathbb{E}\left[V_{0}\right]}{\epsilon}+\frac{\sigma^{2}}{\beta\mu^{2}n\epsilon}$ (13)
$+\frac{\sqrt{L\bar{\rho}_{W}\sigma^{2}}}{\mu^{3/2}\left(1-\bar{\rho}_{W}\right)^{3/2}\epsilon^{1/2}}\Biggl).$
@@
$\hat{K}=\mathcal{O}\left(\frac{L}{\left(1-\bar{\rho}_{W}\right)^{2}\epsilon}+\frac{\sigma^{2}}{\beta n\epsilon^{2}}+\frac{\sqrt{L\bar{\rho}_{W}\sigma^{2}}}{\left(1-\bar{\rho}_{W}\right)^{3/2}\epsilon^{3/2}}\right)\mathbb{E}\left[V_{0}\right].$ (14)
@@
$\hat{K}=\mathcal{O}\Biggl($ $\frac{L}{\left(1-\bar{\rho}_{W}\right)^{2}\epsilon}+\frac{R_{0}}{\epsilon}+\frac{L\sigma^{2}}{\beta n\epsilon^{2}}$ (15)
$+\frac{L\sqrt{\bar{\rho}_{W}\sigma^{2}}}{\left(1-\bar{\rho}_{W}\right)^{3/2}\epsilon^{3/2}}\Biggl)\left(\mathbb{E}\left[f\left(\bar{x}_{0}\right)-f^{*}\right]\right),$
@@
Proof 3.2.
See Section 4.3
Based on the iteration complexity derived in Corollary 1 and the number of local updates and communications per iteration, we can derive the communication and computational complexity of FlexGT for different objective functions, as summarized in Table 2. To the best of our knowledge, FlexGT achieves the best-known complexity among GT-based methods for solving Problem (1) in most settings (cf. Remark 1).
Remark 2
Table 2 highlights the impact of $\alpha$ and $\beta$ on the communication and computational complexity of FlexGT. Unlike previous works [15, 28], our results reveal an inherent trade-off between communication and computation. Particularly, as the number of computation steps $\beta$ increases, the communication complexity decreases while the computational complexity increases. As the number of communication steps $\alpha$ increases, the computational complexity decreases, while the communication complexity may either increase or decrease depending on $\rho_{W}$. This highlights the need for a careful design of $\alpha$ and $\beta$ to balance communication and computational costs according to the specific application requirements.